# Which Reasoning Trajectories Teach Students to Reason Better? A Simple Metric of Informative Alignment

### v1

@@ Abstract
Long chain-of-thought (CoT) trajectories provide rich supervision signals for distilling reasoning from teacher to student LLMs. However, both prior work and our experiments show that trajectories from stronger teachers do not necessarily yield better students, highlighting the importance of data–student suitability in distillation. Existing methods assess suitability primarily through student likelihood, favoring trajectories that closely align with the model’s current behavior but overlooking more informative ones. Addressing this, we propose Rank-Surprisal Ratio (RSR), a simple metric that captures both alignment and informativeness to assess the suitability of a reasoning trajectory. RSR is motivated by the observation that effective trajectories typically combine low absolute probability with relatively high-ranked tokens under the student model, balancing learning signal strength and behavioral alignment. Concretely, RSR is defined as the ratio of a trajectory’s average token-wise rank to its average negative log-likelihood, and is straightforward to compute and interpret. Across five student models and reasoning trajectories from 11 diverse teachers, RSR strongly correlates with post-training performance (average Spearman 0.86), outperforming existing metrics. We further demonstrate its practical utility in both trajectory selection and teacher selection.
$\dagger$ Corresponding authors. Inquiries may be sent to: yumingyang23@m.fudan.edu.cn, zouyicheng@pjlab.org.cn, qz@fudan.edu.cn* Code is available at https://github.com/UmeanNever/RankSurprisalRatio.
@@ 1 Introduction
Yet, stronger reasoning teachers do not necessarily yield better students [22, 11]. Our extensive experiments show that the post-training effectiveness of reasoning trajectories varies substantially across student models, indicating that the suitability between data and student is critical for effective learning. Existing data engineering methods assess data suitability primarily through the student’s probability assignments [47, 18], favoring high-likelihood trajectories that align closely with the model’s current behavior. Such trajectories, however, often provide limited new learning signals. In contrast, more informative trajectories are typically less familiar to the student and thus overlooked by these methods. This leads to a fundamental Informative Alignment challenge: how to identify reasoning data that are both well aligned with the student and sufficiently informative?
To address this challenge, we propose a simple yet effective metric, Rank-Surprisal Ratio (RSR), which quantifies the suitability of a reasoning trajectory for a given student by jointly capturing alignment and informativeness. Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting absolute unfamiliarity and relative familiarity. Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while still containing behavioral patterns within the student’s prior experience, such that their tokens rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
Figure 1: Illustration of the intuition behind RSR. Suitable reasoning trajectories should balance informativeness and alignment by having low absolute probability but relatively high-ranked tokens under the student model.
@@
  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data–student suitability.
  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
@@ 2.1 Experimental Settings
Our teacher-student pairing study involves two major steps: (1) For each teacher model, we prompt it to generate a long CoT response for each math problem in our 5000-problem set (see § A.1), forming a trajectory dataset specific to that teacher. (2) For each teacher–student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance. All students are pretrained base models. To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher–student pair. Reported results are averaged over these three runs. More implementation details are provided in the appendix.
@@
We use 11 reasoning LLMs spanning 4B to 671B parameters across multiple model families, including GPT-OSS [3], DeepSeek [12], Qwen [39], LLaMA-Nemotron [4], and Phi [1].
@@
We evaluate the reasoning performance of the fine-tuned student models on four standard math benchmarks: AIME’25, AIME’24, AMC’23, and MATH500 [14] using the Acc@4 metric, and report results averaged across all benchmarks.
@@ 2.2 Results
The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model. Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) may likewise be ineffective at improving stronger students (Qwen-3-14B). Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb. Overall, we find no simple teacher–student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
@@ 3.1 Surprisal and Rank
We introduce two different methods to quantify the amount of information carried by a token with respect to a student model. They serve as building blocks of our metric.
@@
A common measure is based on the probability of generating the current token $t_{k}$ given its preceding context $\mathbf{c}_{k}=(t_{1},\ldots,t_{k-1})$ under the student model $\theta$. For numerical stability, probabilities are typically transformed into the log space, and the negative log-likelihood is used as a measure of informativeness, also known as Surprisal [13]
@@
Unlike surprisal, rank captures relative familiarity of the token by comparing target token against alternative candidates, revealing signals overlooked by probability-based measures. For instance, a token may receive low absolute probability while still being among the top-ranked candidates.
@@ 3.2 Limitations of Probability-Based Metrics
Existing work primarily relies on surprisal or log-probability to assess data suitability. For example, Zhang et al. [47] selects trajectories based on the average log-probability of response tokens under the student model. Since surprisal is the negation of log-probability, we implement this metric as the average surprisal, denoted as Avg-Surprisal. More recently, Just et al. [18] computes token-level log-probability based on a local context $\mathbf{c}_{k}^{\text{local}}$ (several preceding sentences), which we implement as average local surprisal (Avg-Surplocal). Under these metrics, trajectories with lower surprisal are considered more suitable.
@@ 3.3 Insight and Simulation
The above analysis suggests that effective teacher trajectories should balance the informativeness of the data with their alignment to the student’s current behavior: they should be neither overly similar to the student’s own generations nor excessively deviant from its predicted distribution.
At first glance, this balance may appear to pose a dilemma. However, we argue that it can be resolved by viewing informativeness and alignment through the lens of absolute unfamiliarity and relative familiarity. Informativeness does not require trajectories to be entirely unfamiliar; rather, it suffices that they deviate from the dominant patterns and thus have low absolute probability of being generated by the student. Conversely, alignment does not require exact agreement with the student’s outputs, but instead that the corresponding tokens have relatively higher likelihoods than other candidates in the vocabulary.
Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while still exhibiting behavioral patterns within the student’s prior experience. Therefore, tokens in such trajectories are assigned low absolute probability (high surprisal) by the student model while still ranking relatively high (low rank values) in its prediction distribution.
@@
We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we consider effective. Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$. The first mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant patterns, which arise from student’s massive general training data. The second mode $Z_{B}$, represents tokens whose patterns resemble strong reasoning trajectories, reflecting the student’s prior exposure to relevant data. We instantiate both $Z_{A}$ and $Z_{B}$ as Zipf distributions [52, 26] over $\mathcal{V}$. Then the student’s overall token-level prediction distribution $Z$ is constructed as a mixture of $Z_{A}$ and $Z_{B}$:
@@
Based on $Z$, we simulate four types of reasoning trajectories and examine their token-level surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior. More simulation details are provided in § A.6.
@@
Table 3: Simulation results for different types of trajectories, reporting average token probability, surprisal, rank, and $_{\text{token}}. RSR$$_{\text{token}}$ shows promise as a reliable metric for identifying suitable reasoning trajectories (§ 3.4).
@@
Figure 2 presents the simulated bimodal distribution $Z$, where effective reasoning trajectories are expected to align with the distribution of $Z_{B}$, and thus resemble $X_{B}$. As shown in Table 3, $X_{B}$ exhibits much higher average surprisal (lower absolute probability) than $X_{A}$ and $X_{D}$, while maintaining relatively low rank values (higher-ranked), consistent with the quantitative patterns implied by our analysis. Crucially, the results suggest a promising direction for measuring data suitability, which we formalize in § 3.4.
@@ 3.4 Proposed Metric: Rank-Surprisal Ratio
Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2). This suggests that tokens with higher surprisal have a greater impact on student learning and should be emphasized when computing the average.
@@
The resulting metric yields a concise form that can be interpreted as the ratio of average rank to average surprisal at the trajectory level.
In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability. Since tokens with excessively large ranks are effectively indistinguishable from the student’s perspective, we clip rank values at a threshold $r_{max}$. Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR), as
@@
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates deviation from dominant patterns and provides informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
@@ 4.1 Main Analysis
For each of the five student models, we compute average trajectory-level suitability (or quality) scores for reasoning datasets generated by eleven teacher models (§ 2.1) using different metrics. We then measure the correlation between these dataset-level scores and the student’s average post-training reasoning performance on the corresponding teacher data. We primarily report Spearman’s correlation coefficient, as Pearson correlation exhibits similar trends. We use a clipping threshold $r_{max}=100$ for our RSR metric in all subsequent experiments. Additional analysis details are provided in § A.7, and complete metric scores are reported in § C.2.
@@
We compare RSR against a diverse set of metrics for evaluating reasoning trajectories. These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal; [18]) and rank-based metrics. We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [43], LLM-judged quality scores, and answer accuracy on verifiable questions. In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE; [29]) and influence scores [16].
@@ 4.2 Ablation Study
The derivation of Rank-Surprisal Ratio involves several design components as well as a hyperparameter $r_{max}$. We conduct an ablation study to examine how these choices affect correlation performance.
As shown in Table 5, removing either rank clipping or the surprisal-weighted averaging substantially degrades the correlation, validating the necessity of both components in our metric. In addition, the “Reduced sample size” setting estimates the dataset-level RSR using only 200 trajectories per teacher instead of the full 5,000. The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$. More ablation results are provided in § C.3.
@@ 5.1 Trajectory Selection
The trajectory selection task aims to identify the most effective reasoning trajectory from a set of candidates for a given problem or prompt. We adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion. This procedure is repeated for all 5,000 training problems and all student models, yielding student-specific 5,000-trajectory teacher datasets for each selection method. We then fine-tune student models on the constructed datasets and evaluate their post-training reasoning performance on standard math benchmarks, consistent with previous experiments. We compare RSR against a random selection baseline and multiple previously discussed metrics. For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the score.

### v2

@@ Abstract
Long chain-of-thought (CoT) trajectories provide rich supervision signals for distilling reasoning from teacher to student LLMs. However, both prior work and our experiments show that trajectories from stronger teachers do not necessarily yield better students, highlighting the importance of data-student suitability in distillation. Existing methods assess suitability primarily through student likelihood, favoring trajectories that align closely with the student model’s current behavior but overlooking more informative ones. Addressing this, we propose Rank-Surprisal Ratio (RSR), a simple metric that captures both alignment and informativeness to assess the suitability of a reasoning trajectory. RSR is motivated by the observation that effective trajectories typically balance learning signal strength and behavioral alignment by combining low absolute probability with relatively high-ranked tokens under the student model. Concretely, RSR is defined as the ratio of a trajectory’s average token-wise rank to its average negative log-likelihood, and is straightforward to compute and interpret. Across five student models and reasoning trajectories from 11 diverse teachers, RSR strongly correlates with post-training reasoning performance (average Spearman 0.86), consistently outperforming existing metrics. We further demonstrate its practical utility in both trajectory selection and teacher selection.
$\dagger$ Corresponding authors. Inquiries may be sent to: yumingyang23@m.fudan.edu.cn, zouyicheng@pjlab.org.cn, qz@fudan.edu.cn* Code and data are available at https://github.com/UmeanNever/RankSurprisalRatio.
@@ 1 Introduction
Yet, stronger reasoning teachers do not necessarily yield better students [23, 12]. Our extensive experiments show that the post-training effectiveness of reasoning trajectories varies substantially across student models, indicating that the suitability between data and student is critical for effective learning. Existing data engineering methods assess data suitability primarily through the student’s probability assignments [49, 19], favoring high-likelihood trajectories that align closely with the model’s current behavior. Such trajectories, however, often provide limited new learning signals. In contrast, more informative trajectories are typically less familiar to the student and thus overlooked by these methods. To facilitate more effective learning, it is crucial to strike a balance between familiarity and informativeness, echoing the psychological concept of the zone of proximal development [36]. This leads to a fundamental Informative Alignment challenge: how to identify reasoning data that are both well aligned with the student and sufficiently informative?
To address this challenge, we propose a simple yet effective metric, Rank-Surprisal Ratio (RSR), which quantifies the suitability of a reasoning trajectory for a given student by jointly capturing alignment and informativeness. Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting both absolute unfamiliarity and relative familiarity. Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while still containing behavioral patterns within the student’s prior experience, such that their tokens rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
Figure 1: Illustration of the intuition behind RSR. Suitable reasoning trajectories should balance informativeness and alignment, having low absolute probability while their tokens remain relatively high-ranked under the student model.
@@
  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data-student suitability.
  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory for a given student model by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
@@ 2.1 Experimental Settings
Our teacher-student pairing study involves two major steps: (1) For each teacher model, we prompt it to generate a long CoT response for each math problem in our 5000-problem set (see § A.1), forming a trajectory dataset specific to that teacher. (2) For each teacher–student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance. All students are pre-trained base models. To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher–student pair. Reported results are averaged over these three runs. More implementation details are provided in Appendix A.
@@
We use 11 reasoning LLMs (§ A.3) spanning 4B to 671B parameters across multiple model families, including DeepSeek [13], GPT-OSS [3], Qwen [41], LLaMA-Nemotron [4], and Phi [1].
@@
We evaluate the reasoning performance of fine-tuned student models on four popular math benchmarks (AIME’25, AIME’24, AMC’23, and MATH500 [15]) using the Acc@4 metric (§ A.4), and report results averaged across all benchmarks.
@@ 2.2 Results
The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model. Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) can likewise be ineffective when training stronger students (Qwen-3-14B). Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb. Overall, we find no simple teacher–student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
@@ 3.1 Surprisal and Rank
We first introduce two fundamental methods for quantifying the amount of information a token carries with respect to a student model. They serve as the building blocks of our metric.
@@
A common measure is based on the probability of generating the current token $t_{k}$ given its preceding context $\mathbf{c}_{k}=(t_{1},\ldots,t_{k-1})$ under the student model $\theta$. For numerical stability, probabilities are typically transformed into log space, and the negative log-likelihood—also known as surprisal—is used as a measure of informativeness [14].
@@
Unlike surprisal, rank captures relative familiarity of the token by comparing target token against alternative candidates, revealing signals overlooked by probability-based measures. For instance, a token may be assigned a low absolute probability while still ranking among the top candidates.
@@ 3.2 Limitations of Probability-Based Metrics
Existing work primarily relies on log-probability or surprisal to assess data suitability. For example, Zhang et al. [49] select trajectories based on the average log-probability of response tokens under the student model. Since surprisal is the negation of log-probability, we implement this metric as the average surprisal, denoted as Avg-Surprisal. More recently, Just et al. [19] compute token-level log-probability based on a local context $\mathbf{c}_{k}^{\text{local}}$ (several preceding sentences), which we implement as average local surprisal (Avg-Surplocal, § B.6). Under these metrics, trajectories with lower surprisal are considered more suitable for the student model.
@@ 3.3 Insight and Simulation
The above analysis suggests that suitable (i.e., effective) teacher trajectories should strike a balance between data informativeness and alignment with the student’s current behavior: they should be neither overly similar to the student’s own generations nor excessively deviant from its prediction distribution.
At first glance, this balance may appear to pose a dilemma. However, we argue that it can be resolved by viewing informativeness and alignment through the lens of absolute unfamiliarity and relative familiarity. Informativeness does not require trajectories to be entirely unfamiliar; rather, it suffices that they deviate from the dominant patterns and thus have low absolute probability of being generated by the student. Conversely, alignment does not entail exact agreement with the student’s outputs, but instead requires that the corresponding tokens have relatively higher likelihoods than other candidates in the vocabulary.
Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while still exhibiting behavioral patterns within the student’s prior experience. Therefore, tokens in such trajectories are assigned low absolute probability (i.e., high surprisal) by the student model while still ranking relatively high (i.e., having low rank values) in its prediction distribution.
@@
We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we deem effective. Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$. The first mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant patterns, which arise from student’s massive general training data. The second mode $Z_{B}$, represents tokens whose patterns resemble strong reasoning trajectories, reflecting the student’s prior exposure to relevant data. We instantiate both $Z_{A}$ and $Z_{B}$ as Zipf distributions [54, 27] over $\mathcal{V}$. Then the student’s overall token-level prediction distribution $Z$ is constructed as a mixture of $Z_{A}$ and $Z_{B}$:
@@
Based on $Z$, we simulate four types of reasoning trajectories and examine their average token surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior. More simulation details are provided in § A.6.
@@
[…]
@@
Figure 2 presents the simulated bimodal distribution $Z$, where effective reasoning trajectories are expected to align with the distribution of $Z_{B}$, and thus resemble $X_{B}$. As shown in Table 3, $X_{B}$ exhibits a much higher average surprisal (lower probability) than $X_{A}$ and $X_{D}$, while maintaining relatively low rank values (top-ranked), consistent with the quantitative patterns implied by our prior analysis. Crucially, the results suggest a promising direction for measuring data suitability, which we formalize in § 3.4.
@@ 3.4 Proposed Metric: Rank-Surprisal Ratio
Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2). This suggests that tokens with higher surprisal have a greater impact on student learning and should therefore be emphasized when computing the average.
@@
The resulting metric yields a concise form that can be interpreted as the ratio of the average token rank to the average token surprisal over a trajectory.
In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability. Since tokens with excessively large ranks are effectively indistinguishable from the student’s perspective, we clip rank values at a threshold $r_{max}$. Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR [^2] By default, RSR refers to the trajectory-level RSR in this paper, unless otherwise specified (e.g., in correlation analysis).), as
@@
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates greater deviation from dominant patterns and provides more informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
@@ 4.1 Main Analysis
For each of the five student models and each metric, we aggregate trajectory-level suitability (or quality) scores to obtain dataset-level scores for reasoning datasets generated by eleven teacher models (§ 2.1). We then measure the correlation between these dataset-level scores and the student’s reasoning performance after training on the corresponding teacher-generated datasets. We primarily report Spearman’s correlation coefficient, while Pearson correlation exhibits similar trends (§ C.2). For dataset-level RSR, we adopt a weighted averaging scheme (§ A.8) similar to Eq. 6, which yields slightly higher correlation than a simple average of trajectory-level RSR. We use a clipping threshold of $r_{max}=100$ for RSR in all subsequent experiments. Additional analysis details are provided in § A.7, and complete metric scores are reported in § C.2.
@@
We compare RSR against a diverse set of metrics for evaluating reasoning trajectories. These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal [19]) and rank-based metrics. We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [45], LLM-judged quality scores, and answer accuracy on verifiable questions. In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE [30]) and influence scores [17].
@@ 4.2 Ablation Study
The derivation of Rank-Surprisal Ratio involves several design components, as well as a hyperparameter $r_{max}$. We conduct an ablation study to examine how these choices affect the correlation strength of dataset-level RSR.
As shown in Table 5, removing either rank clipping or the surprisal-weighted averaging substantially degrades the correlation, validating the necessity of both components in our metric. In addition, the “Reduced sample size” setting estimates the dataset-level RSR using only 200 trajectories per teacher instead of the full 5,000. The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$. Additional ablation results are in § C.3.
@@ 5.1 Trajectory Selection
The trajectory selection task aims to identify the most effective reasoning trajectory from a set of candidates for a given problem or prompt. We adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion. This procedure is repeated for all 5,000 training problems and all student models, yielding student-specific 5,000-trajectory teacher datasets for each selection method. We then fine-tune student models on the constructed datasets and evaluate their post-training reasoning performance on standard math benchmarks, consistent with previous experiments. We compare RSR against a random selection baseline and several previously discussed metrics. For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the corresponding trajectory-level score.